OpenStudy (anonymous):

Which of the following is the equation of the axis of symmetry of the quadratic equation y = 2x2 + 24x + 62?

sam (.sam.):

complete the square, y=2x^2 +24x+62 y=2(x^2+12x)+62 y=2(x+6)^2 -36(2)+62 y=2(x+6)^2 -10 axis of symmetry is x=-6

OpenStudy (anonymous):